# Leaking Outside the Box: Kinetic Turbulence with Cosmic-Ray Escape

### v2

@@ Abstract
However, most kinetic simulations to date face a significant shortcoming: they lack a genuine steady state, due to the fact that the simulation domains employed are of finite size and closed, which leads to a progressive accumulation of energy over time and to the pile-up of particles at the highest energies [36]. This presents challenges for connecting simulations and realistic astrophysical objects, which may convert a significant fraction of their available power into escaping cosmic rays. Moreover, it is known that power-law distributions emerge most naturally from a competition between particle acceleration and escape. Therefore, steady-state simulations, if achievable, could offer a convenient and sorely needed testbed for theoretical models.
@@
Employing for simplicity an electron–positron composition ($m_{e^{-}}=m_{e^{+}}=m$) of mean number density $n_{0}$, we explore the effect of different cold plasma magnetizations $\sigma_{0}=B_{0}^{2}/(4\pi n_{0}mc^{2})$ on particle acceleration and escape. Several runs with different $\sigma_{0}=\{0.2,1,4,10,20,80\}$ are performed. A periodic box of size $L^{3}$ is initially filled with thermal particles at temperature $T_{0}$. Their initial displacements $\delta x,\delta y$, which determine the escape condition, are uniformly distributed between $-l_{\rm esc}$ and $l_{\rm esc}$. We set $T_{0}=mc^{2}[\sigma_{0}(2+\sigma_{0})]/[3+3\sigma_{0}]$, which ensures that the mean energy of the injected particles $\mathrm{E}_{0}\sim mc^{2}\sigma_{0}$ is close to the expected steady-state mean particle energy (see Eq. (1) below). To sustain turbulence, a Langevin antenna drives an external current [38] at the largest (box-scale) wavelengths with frequency $\omega_{\rm ant}\approx 0.9\omega_{\rm A}$, slightly off-resonant to the lowest Alfvén frequency of the box $\omega_{\rm A}=(2\pi/L)v_{\rm A}$. The Alfvén speed $v_{\rm A}=c\sqrt{\sigma/(\sigma+1)}$ is defined here based on the initial hot magnetization $\sigma=B_{0}^{2}/(4\pi w_{0}n_{0}mc^{2})=\sigma_{0}/w_{0}$, where $w_{0}=\mathcal{K}_{3}(mc^{2}/T_{0})/\mathcal{K}_{2}(mc^{2}/T_{0})$ is the initial specific enthalpy and $\mathcal{K}_{n}(z)$ are $n$-th order modified Bessel functions of the second kind. The decorrelation rate of the antenna is $\gamma_{\rm ant}=0.5\omega_{\rm ant}$. We focus on the strong-turbulence regime, where the magnetic perturbation driven by the antenna is comparable to the background magnetic field: $\delta b\equiv\delta B/B_{0}\approx 1$. The spatial resolution of all simulations presented in this Letter is $960^{3}$, and the physical box size $L=480d$, where $d=[(1+2\sigma_{0})mc^{2}/(4\pi n_{0}e^{2})]^{1/2}$ is a reference hot-plasma skin depth based on the expected steady-state kinetic temperature (discussed below). The simulations are run for at least 20 light-crossing times $t_{\rm cross}=l_{\rm esc}/c$.
@@
Energy Balance in Steady State.— Our simulations presented in the following show that large-amplitude kinetic turbulence with particle escape settles into a state where the average kinetic and magnetic pressures are of the same order. A supporting analytical argument can be provided by considering the balance between turbulent energization and energy losses from particle escape: $(\mathrm{E}_{\rm esc}-\mathrm{E}_{0})n_{0}/t_{\rm esc}\simeq\delta b^{2}B_{0}^{2}/(4\pi t_{0})$, where $\mathrm{E}_{\rm esc}$ is the mean energy of escaping particles, $t_{\rm esc}$ is the escape time, and $t_{0}$ is the turbulence cascade time. Assuming for simplicity that particles from the thermal population stream along tangled magnetic-field lines, thereby diffusing perpendicularly to the mean magnetic field, we estimate their escape time as $t_{\rm esc}\simeq N\delta t\simeq l_{\rm esc}^{2}/(l_{\parallel}\delta b^{2}v_{\rm th})$, where $\delta t\simeq l_{\parallel}/v_{\rm th}$ is the time to travel one field-parallel correlation length $l_{\parallel}$ at thermal speed $v_{\rm th}\equiv[T/(wm)]^{1/2}$ and $N\simeq[l_{\rm esc}/(l_{\parallel}\delta b)]^{2}$ is the number of correlation lengths needed to diffuse over a distance $l_{\rm esc}$. The cascade time $t_{0}\simeq\chi^{-2}l_{\parallel}/v_{\rm A}\simeq l_{0}^{2}/(l_{\parallel}\delta b^{2}v_{\rm A})$, where $\chi=\delta b\,l_{\parallel}/l_{0}$ is the turbulence nonlinearity parameter and $l_{0}$ is the perpendicular coherence scale. Focusing on the regime of interest with $\sigma_{0}\gtrsim 1$, $\mathrm{E}_{\rm esc}/\mathrm{E}_{0}\gtrsim 1$, and assuming at most mildly relativistic Alfvén speeds, such that $v_{\rm A}^{2}\simeq c^{2}\sigma_{0}/w$ (to be justified below), we finally obtain
@@
where $\beta\simeq 2v_{\rm th}^{2}/v_{\rm A}^{2}\simeq 2T/(\sigma_{0}mc^{2})$ is the plasma beta, i.e., the ratio of plasma kinetic to magnetic pressures. Thus, large-amplitude ($\delta b\sim 1$) magnetized turbulence, driven on scales comparable to the particle escape distance ($l_{0}\sim l_{\rm esc}$), naturally settles into a state with $\beta\sim 1$. This puts a ceiling on the hot plasma magnetization $\sigma\lesssim 1$, thereby limiting the turbulent motions $\delta v\simeq\delta bv_{\rm A}$ to the mildly relativistic regime. In turn, this limits the maximum particle scattering rate and the resulting acceleration efficiency, as shown below.
@@
In Fig. 1(c), we track the evolution of the Larmor radius $\rho_{\rm max}=(\gamma_{\rm max}^{2}-1)^{1/2}mc^{2}/(qB_{0})$ of the highest-energy particles in the box 111we determine $\gamma_{\rm max}={\rm E}_{\rm max}/(mc^{2})+1$ based on the highest energy at which $f({\rm E}_{\max})$ is less than a factor of $10^{5}$ below the peak of $f({\rm E})$. At early times, $\rho_{\rm max}(t)$ grows at near-exponential rate [26], which is followed by saturation at the Hillas limit $\rho_{\rm max}\approx\rho_{\rm Hillas}\approx l_{\rm esc}$. The energization of the plasma and its approach toward steady state is evident also from the evolution of the plasma-beta parameter $\beta\equiv 8\pi n_{0}T/B_{\rm rms}^{2}$ (Fig. 1(d)), where $T=[mc^{2}/(3n_{0})]\int(\gamma-1/\gamma)f(\rm E){\rm dE}$ and $\gamma\equiv{\rm E}/(mc^{2})+1$  [^2] For simplicity, and considering the absence of strong bulk flows, we compute the kinetic and magnetic pressures in the simulation frame, rather than in the proper fluid frame [21]. We confirmed that a more elaborate estimate based on the transformation into the fluid frame leaves results practically unchanged.. From panel (d) of Fig. 1, we can conclude that the average $\beta\approx 1.5$ in steady state. This is a generic feature of all of our runs: for any value of the cold magnetization $\sigma_{0}$, the system eventually relaxes toward a state in which the plasma kinetic and magnetic pressures equilibrate, in line with the theoretical argument presented above. Such a universal behavior at high magnetization introduces an upper limit on the bulk turbulent motions of the plasma. Since the hot magnetization $\sigma=\sigma_{0}/w\approx 0.33$ (where we used $w\approx 4T/(mc^{2})\approx 3\sigma_{0}$), the typical Alfvén speed is $v_{\rm A}\approx 0.5c$, implying that the bulk motions are at most mildly relativistic, at least in the absence of radiative-cooling mechanisms [16, 22, 25].
@@
Figure 2: (a) Steady-state particle distributions of the contained population. Stars denote the Hillas limit. The inset shows the measured power-law indices for both contained (green line) and escaping (pink line) particle populations. (b) The number fraction of nonthermal particles (solid lines) and the fraction of energy attributed to nonthermal particles (dashed lines) for the contained (green) and escaping (pink) population. (c) The measured steady-state plasma $\beta$ for different cold magnetizations $\sigma_{0}$.
@@
How the plasma magnetization $\sigma_{0}$ shapes the nonthermal particle distributions is shown in Fig. 2. Fig. 2(a) shows the steady-state particle distributions as a function of ${\rm E}/(\sigma_{0}mc^{2})$. Due to the linear dependence $T\propto\sigma_{0}mc^{2}$ [see Eq. (1)], the distribution peaks approximately overlap, albeit with a slight shift toward lower energies for higher $\sigma_{0}$. Efficient acceleration is enabled in all cases considered, in the sense that all distributions form extended nonthermal tails reaching the maximum energy set by the system-size (Hillas) limit ($\rho_{\rm max}\approx l_{\rm esc}$; marked with star symbols in panel (a)). More strongly magnetized systems feature harder nonthermal tails, as shown in the inset of Fig. 2(a). The obtained power-law slopes are broadly consistent with previous kinetic simulations of strong turbulence in closed domains [15, 27, 19, 22, 23].
@@
In strongly magnetized turbulence, particles can be injected into the nonthermal population via nonideal electric fields at reconnecting current sheets [19]. Following the reconnection scenario, we can estimate the fraction of nonthermal particles $\xi_{\rm nt}$ from the expected rate of reconnection in relativistic plasmas [43]. In one turbulence eddy turnover time, reconnecting sheets process a fraction ${\mathcal{V}}_{\rm rec}\sim\beta_{\rm rec}$ of the plasma volume [19], where $\beta_{\rm rec}$ is the reconnection rate in units of $B_{0}v_{\rm A}$, assuming $\delta B\sim B_{0}$ (i.e., moderate-guide-field reconnection). The reconnection rate can be obtained from the slope $S$ of the separatrix field lines (defining the reconnection exhaust) as $\beta_{\rm rec}\simeq S(1-S^{2})/(1+S^{2})$ [44, 45, 43]. For a cold plasma upstream of the reconnection layer  [^3] Strictly speaking, we are implicitly assuming that reconnection layers both heat the plasma and accelerate particles, so that the plasma flowing into the reconnection layers can be modeled as cold in the first approximation., $S$ is given by the following nonlinear equation [43]:
@@
where $\sigma_{\rm m}\approx\sigma_{0}(1-S^{2})/(1+S^{2})$ is the “microscale” magnetization parameter, and we took half of the asymptotic X-line kinetic pressure to obtain the above formula [43]. Assuming that an order-unity fraction of particles reprocessed by reconnection is injected into the acceleration process, we finally estimate the nonthermal fraction as $\xi_{\rm nt}\sim\beta_{\rm rec}$. Our estimate $\xi_{\rm nt}\approx 0.85\beta_{\rm rec}$ (with 0.85 as an ad hoc prefactor) is shown with a black dashed line in Fig. 2(b), demonstrating good agreement with the simulations. This provides further evidence that particle injection into the acceleration process is controlled by magnetic reconnection.
@@
Figure 3: (a) Energy diffusion coefficient versus energy for different $\sigma_{0}$. Dashed black lines indicate fits of the form $D_{\rm EE}=t_{\rm acc}{\rm E}^{2}$, where $t_{\rm acc}$ is the acceleration time. (b) Escape time $t_{\rm esc}$ versus energy for different $\sigma_{0}$. The inset shows the acceleration time $t_{\rm acc}(\sigma_{0})$, which is well-described by a fit of the form $t_{\rm acc}c/l_{\rm esc}\approx 3/\sigma_{0}+2.8$.
@@
In Fig. 3, we study the statistics of particle acceleration and escape. We obtain the acceleration time scale $t_{\rm acc}$ by measuring the energy diffusion coefficient $D_{\rm EE}(\rm E)$ [19, 20], which defines the acceleration time as $t_{\rm acc}(\rm E)=\rm{E}^{2}/D_{\rm EE}(\rm E)$. To this end, we track simulation particles located in different energy bins starting from a time $t_{*}$  [^4] The particle energies are computed in the frame comoving with the drift velocity $c\mbox{\boldmath{$E$}}\times\mbox{\boldmath{$B$}}/B^{2}$ and averaged over the gyration time, to remove high-frequency oscillations., and we obtain the energy-diffusion coefficient as $D_{\rm EE}({\rm E})=\langle\Delta{\rm E}(t)^{2}\rangle/\Delta t$, where $\Delta{\rm E}(t)={\rm E}(t)-{\rm E}(t_{*})$, $\Delta t=t-t_{*}\approx l_{\rm esc}/c$, and the angular brackets represent an average over all particles in a given energy bin. The dependence of diffusion coefficients on energy for different magnetizations is shown in Fig. 3(a). The diffusion coefficients scale as $D_{\rm EE}\propto{\rm E}^{2}$, consistent with previous measurements performed in closed domains [19, 20]. A slight departure from the $\propto{\rm E}^{2}$ scaling is seen at high energies, approaching the Hillas limit. By performing fits of the form $D_{\rm EE}(\rm E)=t_{\rm acc}E^{2}$ (dashed black lines), we determine the acceleration time $t_{\rm acc}(\sigma_{0})$ (inset of Fig. 3(b)). The efficiency of acceleration is related to the speed of the scattering agents [34], which typically move with the velocity of the turbulent bulk flow $\delta v\sim v_{\rm A}$. The rate of stochastic acceleration is then limited by the largest achievable steady-state bulk velocity. This is reflected in our fitting formula $t_{\rm acc}\approx(3/\sigma_{0}+2.8)l_{\rm esc}/c$. For $\sigma_{0}\gg 1$, the acceleration time settles at $t_{\rm acc}\approx 2.8l_{\rm esc}/c$, which represents the asymptotic value set by the heating-regulated ceiling for the Alfvén speed.

### v3

@@ Abstract
However, most kinetic simulations to date face a significant shortcoming: they lack a genuine steady state, due to the fact that the simulation domains employed are of finite size and closed, which leads to a progressive accumulation of energy over time and to the pile-up of particles at the highest energies [17, 36, 37]. This presents challenges for connecting simulations and realistic astrophysical objects, which may convert a significant fraction of their available power into escaping cosmic rays. Moreover, it is known that power-law distributions emerge most naturally from a competition between particle acceleration and escape. In closed domains, the local trapping of particles provides an effective “escape” mechanism [36], enabling the development of power-law distributions. However, without an energy sink to balance the turbulent forcing, such power-law distributions are only transient states, followed by progressive energy pile-up. Therefore, steady-state simulations, if achievable, could offer a convenient and sorely needed testbed for theoretical models.
@@
Employing for simplicity an electron–positron composition ($m_{e^{-}}=m_{e^{+}}=m$) of mean number density $n_{0}$, we explore the effect of different cold plasma magnetizations $\sigma_{0}=B_{0}^{2}/(4\pi n_{0}mc^{2})$ on particle acceleration and escape. Several runs with different $\sigma_{0}=\{0.2,1,4,10,20,80\}$ are performed. A periodic box of size $L^{3}$ is initially filled with thermal particles at temperature $T_{0}$. We set $T_{0}=mc^{2}[\sigma_{0}(2+\sigma_{0})]/[3+3\sigma_{0}]$, which ensures that the mean energy of the injected particles $\mathrm{E}_{0}\sim mc^{2}\sigma_{0}$ is close to the expected steady-state mean particle energy (see Eq. (1) below). To sustain turbulence, a Langevin antenna drives an external current [40] at the largest (box-scale) wavelengths with frequency $\omega_{\rm ant}\approx 0.9\omega_{\rm A}$, slightly off-resonant to the lowest Alfvén frequency of the box $\omega_{\rm A}=(2\pi/L)v_{\rm A}$. The Alfvén speed $v_{\rm A}=c\sqrt{\sigma/(\sigma+1)}$ is defined here based on the initial hot magnetization $\sigma=B_{0}^{2}/(4\pi w_{0}n_{0}mc^{2})=\sigma_{0}/w_{0}$, where $w_{0}=\mathcal{K}_{3}(mc^{2}/T_{0})/\mathcal{K}_{2}(mc^{2}/T_{0})$ is the initial specific enthalpy and $\mathcal{K}_{n}(z)$ are $n$-th order modified Bessel functions of the second kind. The decorrelation rate of the antenna is $\gamma_{\rm ant}=0.5\omega_{\rm ant}$. We focus on the strong-turbulence regime, where the magnetic perturbation driven by the antenna is comparable to the background magnetic field: $\delta b\equiv\delta B/B_{0}\approx 1$. The spatial resolution of all simulations presented in this Letter is $960^{3}$, and the physical box size $L=480d$, where $d=[(1+2\sigma_{0})mc^{2}/(4\pi n_{0}e^{2})]^{1/2}$ is a reference hot-plasma skin depth based on the expected steady-state kinetic temperature (discussed below). We use 16 particles per cell in all runs  [^2] For a convergence check in particle number, see e.g. [19].. The simulations are run for at least 20 light-crossing times $t_{\rm cross}=l_{\rm esc}/c$.
@@
Energy Balance in Steady State.— Our simulations presented in the following show that large-amplitude kinetic turbulence with particle escape settles into a state where the average kinetic and magnetic pressures are of the same order. A supporting analytical argument can be provided by considering the balance between turbulent energization and energy losses from particle escape: $(\mathrm{E}_{\rm esc}-\mathrm{E}_{0})n_{0}/t_{\rm esc}\simeq\delta b^{2}B_{0}^{2}/(4\pi t_{0})$, where $\mathrm{E}_{\rm esc}$ is the mean energy of escaping particles, $t_{\rm esc}$ is the escape time, and $t_{0}$ is the turbulence cascade time. Assuming for simplicity that particles from the thermal population stream along tangled magnetic-field lines, thereby diffusing perpendicularly to the mean magnetic field, we estimate their escape time as $t_{\rm esc}\simeq N\delta t\simeq l_{\rm esc}^{2}/(l_{\parallel}\delta b^{2}v_{\rm th})$, where $\delta t\simeq l_{\parallel}/v_{\rm th}$ is the time to travel one field-parallel correlation length $l_{\parallel}$ at thermal speed $v_{\rm th}\equiv[T/(wm)]^{1/2}$ and $N\simeq[l_{\rm esc}/(l_{\parallel}\delta b)]^{2}$ is the number of correlation lengths needed to diffuse over a distance $l_{\rm esc}$. The cascade time $t_{0}\simeq\chi^{-2}l_{\parallel}/v_{\rm A}\simeq l_{0}^{2}/(l_{\parallel}\delta b^{2}v_{\rm A})$, where $\chi=\delta b\,l_{\parallel}/l_{0}$ is the turbulence nonlinearity parameter and $l_{0}$ is the perpendicular coherence scale. Focusing on the regime of interest with $\sigma_{0}\gtrsim 1$, $(\mathrm{E}_{\rm esc}-\mathrm{E}_{0})/\mathrm{E}_{0}\gtrsim 1$, and assuming at most mildly relativistic Alfvén speeds, such that $v_{\rm A}^{2}\simeq c^{2}\sigma_{0}/w$ (to be justified below), we finally obtain
@@
where $\beta\simeq 2v_{\rm th}^{2}/v_{\rm A}^{2}\simeq 2T/(\sigma_{0}mc^{2})$ is the plasma beta, i.e., the ratio of plasma kinetic to magnetic pressures. Thus, large-amplitude ($\delta b\sim 1$) magnetized turbulence, driven on scales comparable to the particle escape distance ($l_{0}\sim l_{\rm esc}$), naturally settles into a state with $\beta\sim 1$. This enforces an upper limit on the hot plasma magnetization $\sigma\lesssim 1$, thereby limiting the turbulent motions $\delta v\simeq\delta bv_{\rm A}$ to the mildly relativistic regime. In turn, this limits the maximum particle scattering rate and the resulting acceleration efficiency, as shown below.
@@
In Fig. 1(c), we track the evolution of the Larmor radius $\rho_{\rm max}=(\gamma_{\rm max}^{2}-1)^{1/2}mc^{2}/(qB_{0})$ of the highest-energy particles in the box 333we determine $\gamma_{\rm max}={\rm E}_{\rm max}/(mc^{2})+1$ based on the highest energy at which ${\rm E}_{\max}f({\rm E}_{\max})$ is less than a factor of $10^{5}$ below the peak of ${\rm E}f({\rm E})$.. As appropriate for $\delta b\sim 1$, we assume that the highest-energy particles have large pitch angles [19]; for $\delta b\ll 1$ the (small) pitch angle should be included in the estimate of $\rho_{\rm max}$ [45, 46]. At early times, $\rho_{\rm max}(t)$ in Fig. 1(c) grows at near-exponential rate [26], which is followed by saturation at the Hillas limit $\rho_{\rm max}\approx\rho_{\rm Hillas}\approx l_{\rm esc}$. The energization of the plasma and its approach toward steady state is evident also from the evolution of the plasma-beta parameter $\beta\equiv 8\pi n_{0}T/B_{\rm rms}^{2}$ (Fig. 1(d)), where $T=[mc^{2}/(3n_{0})]\int(\gamma-1/\gamma)f(\rm E){\rm dE}$ and $\gamma\equiv{\rm E}/(mc^{2})+1$  [^4] For simplicity, and considering the absence of strong bulk flows, we compute the kinetic and magnetic pressures in the simulation frame, rather than in the proper fluid frame [21]. We confirmed that a more elaborate estimate based on the transformation into the fluid frame leaves results practically unchanged.. From panel (d) of Fig. 1, we can conclude that the average $\beta\approx 1.5$ in steady state. This is a generic feature of all of our runs: for any value of the cold magnetization $\sigma_{0}$, the system eventually relaxes toward a state in which the plasma kinetic and magnetic pressures equilibrate, in line with the theoretical argument presented above. Such a universal behavior at high magnetization introduces an upper limit on the bulk turbulent motions of the plasma. Since the hot magnetization $\sigma=\sigma_{0}/w\approx 0.33$ (where we used $w\approx 4T/(mc^{2})\approx 3\sigma_{0}$), the typical Alfvén speed is $v_{\rm A}\approx 0.5c$, implying that the bulk motions are at most mildly relativistic, at least in the absence of radiative-cooling mechanisms [16, 22, 25].
@@
Figure 2: (a) Steady-state particle distributions of the contained population. Stars denote the Hillas limit. The inset shows the measured power-law indices for both contained (green line) and escaping (pink line) particle populations. (b) The number fraction of nonthermal particles (solid lines) and the fraction of energy attributed to nonthermal particles (dashed lines) for the contained (green) and escaping (pink) population. (c) The measured steady-state plasma $\beta$ for different cold magnetizations $\sigma_{0}$. For simulations with $\delta b<1$ see Appendix A.
@@
How the plasma magnetization $\sigma_{0}$ shapes the nonthermal particle distributions is shown in Fig. 2. Fig. 2(a) shows the steady-state particle distributions as a function of ${\rm E}/(\sigma_{0}mc^{2})$. Due to the linear dependence $T\propto\sigma_{0}mc^{2}$ (see Eq. (1)), the distribution peaks approximately overlap, albeit with a slight shift toward lower energies for higher $\sigma_{0}$. Efficient acceleration is enabled in all cases considered, in the sense that all distributions form extended nonthermal tails reaching the maximum energy set by the system-size (Hillas) limit ($\rho_{\rm max}\approx l_{\rm esc}$; marked with star symbols in panel (a)). More strongly magnetized systems feature harder nonthermal tails, as shown in the inset of Fig. 2(a). The obtained power-law slopes are broadly consistent with previous kinetic simulations of strong turbulence in closed domains [15, 27, 19, 22, 23].
@@
In strongly magnetized turbulence, particles can be injected into the nonthermal population at reconnecting current sheets [19]. Following the reconnection scenario, we can estimate the fraction of nonthermal particles $\xi_{\rm nt}$ from the expected rate of reconnection in relativistic plasmas [48]. In one turbulence eddy turnover time, reconnecting sheets process a fraction ${\mathcal{V}}_{\rm rec}\sim\beta_{\rm rec}$ of the plasma volume [19], where $\beta_{\rm rec}$ is the reconnection rate in units of $B_{0}v_{\rm A}$, assuming $\delta B\sim B_{0}$ (i.e., moderate-guide-field reconnection). The reconnection rate can be obtained from the slope $S$ of the separatrix field lines (defining the reconnection exhaust) as $\beta_{\rm rec}\simeq S(1-S^{2})/(1+S^{2})$ [49, 50, 48]. Following [48], we express 555 Eq. (2) is formally derived assuming a cold plasma upstream of the reconnection layer [48]. We adopt it as a crude estimate, given that a more precise expression is not known to us.:
@@
where $\sigma_{\rm m}\approx\sigma_{0}(1-S^{2})/(1+S^{2})$ is the “microscale” magnetization parameter, and we took half of the asymptotic X-line kinetic pressure to obtain the above formula [48]. Assuming that an order-unity fraction of particles reprocessed by reconnection is injected into the acceleration process [52], we finally estimate the nonthermal fraction as $\xi_{\rm nt}\sim\beta_{\rm rec}$. Our estimate $\xi_{\rm nt}\approx 0.85\beta_{\rm rec}$ (with 0.85 as an ad hoc prefactor) is shown with a black dashed line in Fig. 2(b), demonstrating good agreement with the simulations. We note that particles need not be injected exclusively by the nonideal electric fields at reconnecting current sheets [53]; different forms of injection are still debated (e.g., [28, 54]). In particular, with decreasing $\sigma_{0}$ the effect of nonideal fields is expected to diminish, as the particles become less magnetized.
@@
Figure 3: (a) Energy diffusion coefficient versus energy for different $\sigma_{0}$. Dashed black lines indicate fits of the form $D_{\rm EE}={\rm E}^{2}/t_{\rm acc}$, where $t_{\rm acc}$ is the acceleration time. (b) Escape time $t_{\rm esc}$ versus energy for different $\sigma_{0}$. The inset shows the acceleration time $t_{\rm acc}(\sigma_{0})$, which is well-described by a fit of the form $t_{\rm acc}c/l_{\rm esc}\approx 3/\sigma_{0}+2.8$.
@@
In Fig. 3, we study the statistics of particle acceleration and escape. We obtain the acceleration time scale $t_{\rm acc}$ by measuring the energy diffusion coefficient $D_{\rm EE}(\rm E)$ [19, 20], which defines the acceleration time as $t_{\rm acc}(\rm E)=\rm{E}^{2}/D_{\rm EE}(\rm E)$. To this end, we track simulation particles located in different energy bins starting from a time $t_{*}$  [^6] The particle energies are computed in the frame comoving with the drift velocity $c\mbox{\boldmath{$E$}}\times\mbox{\boldmath{$B$}}/B^{2}$ and averaged over the gyration time, to remove high-frequency oscillations., and we obtain the energy-diffusion coefficient as $D_{\rm EE}({\rm E})=\langle\Delta{\rm E}(t)^{2}\rangle/\Delta t$, where $\Delta{\rm E}(t)={\rm E}(t)-{\rm E}(t_{*})$, $\Delta t=t-t_{*}\approx l_{\rm esc}/c$, and the angular brackets represent an average over all particles in a given energy bin. The dependence of diffusion coefficients on energy for different magnetizations is shown in Fig. 3(a). The diffusion coefficients scale as $D_{\rm EE}\propto{\rm E}^{2}$, consistent with previous measurements performed in closed domains [19, 20]. A slight departure from the $\propto{\rm E}^{2}$ scaling is seen at high energies, approaching the Hillas limit. By performing fits of the form $D_{\rm EE}(\rm E)=E^{2}/t_{\rm acc}$ (dashed black lines), we determine the acceleration time $t_{\rm acc}(\sigma_{0})$ (inset of Fig. 3(b)). The efficiency of acceleration is related to the speed of the scattering agents [34], which typically move with the velocity of the turbulent bulk flow $\delta v\sim v_{\rm A}$. The rate of stochastic acceleration is then limited by the largest achievable steady-state bulk velocity. This is reflected in our fitting formula $t_{\rm acc}\approx(3/\sigma_{0}+2.8)l_{\rm esc}/c$. For $\sigma_{0}\gg 1$, the acceleration time settles at $t_{\rm acc}\approx 2.8l_{\rm esc}/c$, which represents the asymptotic value set by the heating-regulated ceiling for the Alfvén speed.